OpenStudy (anonymous):

A rotating beacon is located 15 ft from a wall. If the distance from the beacon to the point on the wall where the beacon is aimed is given by a = 15|sec (2πt)| , where t is in seconds, find a when t = 0.37 seconds. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

\[15*\text{Abs}\left[\text{Sec}\left[2 \pi \frac{37}{100}\right]\right]\text{= }21.91 \text{ ft}\]